OpenStudy (anonymous):

terry worked 12 1/2 hr over a three day period. if she worked 4 1/2 hr the first day and 3 1/4 hr the second, how many hours did terry work on the third day?

OpenStudy (anonymous):

Let x be the hours she worked in the 3rd day Simplify all of them so they all have the same common denominator 25/2=9/2+13/4+x 50/4=18/4+13/4+x 19/4=x